OpenStudy (anonymous):

Giving Medal.. Can Anyone Help me... Been literally half an hour and no answer still

OpenStudy (anonymous):

Simplify: \[\frac{ a ^{-2}b ^{3}c }{ a ^{5}bc ^{2} d ^{0}}\]

OpenStudy (anonymous):

Seriously... Everyone keeps quitting... Can anyone help? .-.

OpenStudy (anonymous):

Wow Open Study is so useless. People only help people if they are asking elementary school questions .-.

OpenStudy (xapproachesinfinity):

that is hard to do lol

OpenStudy (anonymous):

Its Algebra 1 .-. from middle schiol. My highschools doing a reveiw

OpenStudy (xapproachesinfinity):

\(\Huge a^n\times a^m=a^{n+m} \\\Huge \frac{a^n}{a^m}=a^{n-m}\)

OpenStudy (xapproachesinfinity):

that is what you should know and do here to get that done!

OpenStudy (xapproachesinfinity):

what is \(d^0=?\)

OpenStudy (anonymous):

so a^-2 over a^5 = a^3?

OpenStudy (anonymous):

d^0 = 1

OpenStudy (xapproachesinfinity):

the rule said you subtract the power \(\Large a^{-2}/a^5=a^{-2-5}=a^{-7}=\frac{1}{a}\)

OpenStudy (xapproachesinfinity):

i mean 1/a^7

OpenStudy (xapproachesinfinity):

yes that's correct!

OpenStudy (anonymous):

ok let me work this out and see if im right

OpenStudy (xapproachesinfinity):

okay!

OpenStudy (anonymous):

ughh

OpenStudy (xapproachesinfinity):

what is it?

OpenStudy (anonymous):

this is what i have so far .-.